A charged particle moving at a speed of 175 m/s due east, through a magnetic field of 5.00 mT pointing straight up toward the sky, experiences a force of 4.25 ?N directed due north. What is the charge on the particle? Be sure to indicate the sign.
Answer #1

Its a negative charge because the direction of force is opposite to what is given by right hand thumb rule.
Use F =qvB
4.25 *10^(-6)N =q*{175}*{5*10^(-3)}

q={4.25*10^(-3)}/175*5
On solving we get
q=4.85 * 10^(-6)
q= - 4.85 ?C
Answer #2

magnetic force on the charged particle is F = qvB sin theta

so charge q = F/(vB sin theta)

q = 4.25*10^-6/(175* 5*10^-3)

charge q = 4.85 uC it is a +vely charged particle